[figure: 9.7 cm]

Magnon-phonon interaction in (VO)_{2}P_{2}O_{7} studied by Raman scattering

U. Kuhlmann and C. Thomsen^{a}

A.V. Prokofiev, F. Büllesfeld, E. Uhrig, M. Anton, C. Gross and W. Assmus^{b}

aInstitut für Festkörperphysik, Technische Universität Berlin, Hardenbergstr. 36,

D-10623 Berlin, Germany bPhysikalisches Institut, Universität Frankfurt, D-60054 Frankfurt, Germany

We investigated with Raman scattering under high pressure and as a function of temperature the magnetic and phononic excitations in the low-dimensional spin compound (VO)_{2}P_{2}O_{2} (VOPO). In the ambient pressure (AP) phase VOPO comprises two structurally different antiferromagnetic alternating spin chains which are supposed to degenerate to a single chain in the high pressure (HP) phase. In our experiments we studied the interaction of low-energy phonon modes with the spin system of both phases and compared our results with a model of Uhrig et al. (PRB 63 134418 (2001)) developed for the AP phase. We found that the small crystallographic differences between the two phases result in a significant alteration of the magnon-phonon coupling mechanism.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : U. Kuhlmann

Address for correspondence : Institut für Festkörperphysik Sek. PN 5-4 Technische Universität Berlin Hardenbergstr. 36 D-10623 Berlin, Germany

Email address : kuhli@physik.tu-berlin.de